# A Practical Introduction to Kernel Discrepancies: MMD, HSIC & KSD

### v1

@@ 2.3 KSD: Kernel Stein Discrepancy
where $H^{XX}=\big{(}h_{P}(X_{i},X_{j})\big{)}_{1\leq i,j\leq n}$ with the Stein kernel $h_{P}$ as in Equation 43, where $\bar{H}^{XX}$ is the matrix $H^{XX}$ with diagonal entries set to zero, and where $\bm{1}$ is a vector of ones of size $n$. The V-statistic corresponds to the plugin estimator $\mathrm{KSD}_{P}^{2}(\widehat{Q})$ which is strictly positive and can hence be used as an estimator for the (non-squared) KSD by taking its square root, unlike the U-statistic which is unbiased but can be negative. For consistency, the core KSD function $h_{k}^{\operatorname{KSD}}$ can be defined as $h_{P}$ itself.
@@
for all $x\in\mathbb{R}^{d}$ and for ${\bm{f}}=(f_{1},\dots,f_{d})$ where $f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ for $i=1,\dots,d$. This is a Stein operator since $\mathbb{E}_{P}[({\mathcal{A}}_{P}^{\prime}{\bm{f}})(X)]=0$ for all ${\bm{f}}$. Hence, we can define a Stein Discrepancy using this Stein operator, which we kernelise using an RKHS ${\mathcal{H}}$ with reproducing kernel $k$ and unit ball =${\mathcal{F}}=\{{\bm{f}}\in{\mathcal{H}^{d}}:\|{\bm{f}}\|_{{\mathcal{H}^{d}}}\leq 1\}$,, as
@@ 3 Efficient kernel discrepancies estimators
As seen in Equations 11, 21 and 52, MMD, HSIC and KSD can be estimated using one-sample second-order V-statistics, which are estimators of the quantity
@@ 4 Kernel pooling: adaptive kernel discrepancies estimators
As aforementioned, the kernel choices for the MMD, HSIC and KSD estimators crucially affect their utility. To overcome this, we rely on kernel pooling, which consists in combining multiple statistics with different kernels, to construct estimators which are adaptive to the kernel selection. We first explain how the statistics can be normalised to be compared against each other, we then present three kernel pooling methods for combining them, and finally we propose a parameter-free method for constructing a collection of kernels. We note that kernel pooling can be used either with or without normalisation. In practice, we recommend using fuse kernel pooling (Equation 77) with normalisation (Equation 73), which is studied in depth in Biggs et al. (2023).
@@
While this appears to be similar to studentisation, we emphasise that the aim is different: we are not interested in obtaining asymptotic normality but in being able to compare all the normalised statistics $S_{1}/\sigma_{1},\dots,S_{|{\mathcal{K}}|}/\sigma_{|{\mathcal{K}}|}$ in a meaningful way. We note that, for studentisation, there is no real consensus in the literature on which form the estimated standard deviation should take. Here, we propose to use a simple one which aligns well with our study of different types of statistics in Section 3. The unnormalised case simply corresponds to using $\sigma_{k}=1$.
@@
$S_{\overline{k}}=\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\ S_{k}\qquad\textrm{ where }\qquad{\overline{k}}\coloneqq\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}k.$ (75)

### v2

@@ 2.3 KSD: Kernel Stein Discrepancy
[…]
@@
for all $x\in\mathbb{R}^{d}$ and for ${\bm{f}}=(f_{1},\dots,f_{d})$ where $f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ for $i=1,\dots,d$. This is a Stein operator since $\mathbb{E}_{P}[({\mathcal{A}}_{P}^{\prime}{\bm{f}})(X)]=0$ for all $\bm{f}$}. Hence, we can define a Stein Discrepancy using this Stein operator, which we kernelise using an RKHS $\mathcal{H}$} with reproducing kernel $k$ and unit ball ${\mathcal{F}}=\{{\bm{f}}\in{\mathcal{H}^{d}}:\|{\bm{f}}\|_{{\mathcal{H}^{d}}}\leq 1\}$,, as
@@ 3 Efficient kernel discrepancies estimators
As seen in Equations˜11, 21 and 52, MMD, HSIC and KSD can be estimated using one-sample second-order V-statistics, which are estimators of the quantity
@@ 4 Kernel pooling: adaptive kernel discrepancies estimators
As aforementioned, the kernel choices for the MMD, HSIC and KSD estimators crucially affect their utility. To overcome this, we rely on kernel pooling, which consists in combining multiple statistics with different kernels, to construct estimators which are adaptive to the kernel selection. We first explain how the statistics can be normalised to be compared against each other, we then present three kernel pooling methods for combining them, and finally we propose a parameter-free method for constructing a collection of kernels. We note that kernel pooling can be used either with or without normalisation. In practice, we recommend using fuse kernel pooling (Equation˜77) with normalisation (Equation˜73), which is studied in depth in Biggs et al. (2023).
@@
While this appears to be similar to studentisation, we emphasise that the aim is different: we are not interested in obtaining asymptotic normality but in being able to compare all the normalised statistics $S_{1}/\sigma_{1},\dots,S_{|{\mathcal{K}}|}/\sigma_{|{\mathcal{K}}|}$ in a meaningful way. We note that, for studentisation, there is no real consensus in the literature on which form the estimated standard deviation should take. Here, we propose to use a simple one which aligns well with our study of different types of statistics in Section˜3. The unnormalised case simply corresponds to using $\sigma_{k}=1$.
@@
[…]